What is the moment of inertia of a wheel if the mass of the wheel is 5 kg and located 24 inches from the Center of Mass? (ignore mass of spokes and wheel hub)
Answer #1

Moment of inertia of wheel Iw = Mw*L2

1 inch = 0.0254 meters

24 inches = 0.6096 meters

ignoring mass of spokes Is = 0

Iw = 5*0.60962 = 1.85 kg.m2